Pubblicazioni su rivista

[1] S. Aiello, F. Ameli, M. Anghinolfi, et al. Measurement of the atmospheric muon depth intensity relation with the nemo phase-2 tower. Astroparticle Physics, 66:1–7, 2015. cited By 0.

[2] F. Acernese, M. Agathos, K. Agatsuma, et al. Advanced virgo: A secondgeneration interferometric gravitational wave detector. Classical and Quantum Gravity, 32(2), 2015. cited By 1.

[3] J. Aasi, B.P. Abbott, R. Abbott, et al. Narrow-band search of continuous gravitational-wave signals from crab and vela pulsars in virgo vsr4 data. Physical Review D - Particles, Fields, Gravitation and Cosmology, 91(2), 2015. cited By 0.

[4] E. Calloni, M. De Laurentis, R. De Rosa, et al. Towards weighing the condensation energy to ascertain the archimedes force of vacuum. Physical Review D - Particles, Fields, Gravitation and Cosmology, 90(2), 2014. cited By 0.

[5] F. Ambrosino, A. Anastasio, D. Basta, et al. The mu-ray project: Detector technology and first data from mt. vesuvius. Journal of Instrumentation, 9(2), 2014. cited By 0.

[6] S. Adrin-Martnez, M. Ageron, F. Aharonian, et al. Deep sea tests of a prototype of the km3net digital optical module: Km3net collaboration. European Physical Journal C, 74(9), 2014. cited By 0.

[7] T. Accadia, F. Acernese, M. Agathos, et al. Reconstruction of the gravitational wave signal h(t) during the virgo science runs and independent validation with a photon calibrator. Classical and Quantum Gravity, 31(16), 2014. cited By 0.

[8] J. Aasi, B.P. Abbott, R. Abbott, et al. Improved upper limits on the stochastic gravitational-wave background from 2009-2010 ligo and virgo data. Physical Review Letters, 113(23), 2014. cited By 1.

[9] M.G. Aartsen, M. Ackermann, J. Adams, et al. Multimessenger search for sources of gravitational waves and high-energy neutrinos: Initial results for ligo-virgo and icecube. Physical Review D - Particles, Fields, Gravitation and Cosmology, 90(10), 2014. cited By 0.

[10]